# An Empirical Study of LLM-as-a-Judge for LLM Evaluation: Fine-tuned Judge Models are Task-specific Classifiers

### v1

@@ Abstract
Recently, there has been a growing trend of utilizing Large Language Model (LLM) to evaluate the quality of other LLMs. Many studies have employed proprietary close-source models, especially GPT4, as the evaluator. Alternatively, other works have fine-tuned judge models based on open-source LLMs as the evaluator. In this study, we conduct an empirical study of different judge models on their evaluation capability. Our findings indicate that although the fine-tuned judge models achieve high accuracy on in-domain test sets, even surpassing GPT4, they are inherently task-specific classifiers, and their generalizability and fairness severely underperform GPT4.
@@ 1 Introduction
Recently, the evaluation for Large-scale Language Models (LLMs) has drawn considerate attention of research community Liang et al. (2022); Chang et al. (2023). As the capabilities of LLMs continue to develop across various tasks, it is essential to evaluate them from comprehensive perspectives Qin et al. (2023). However, traditional evaluation metrics for generative models, such as BLEU and ROUGE, only capture limited aspects of a model’s performance Mathur et al. (2020).
Figure 1: The general training and inference procedure of fine-tuned judge models.
@@
Some research has proposed LLM-as-a-Judge Li et al. (2023b); Zheng et al. (2023), namley utilizing proprietary LLMs, especially GPT4 Achiam et al. (2023), to evaluate the LLM’s response. By defining evaluation schemes in the prompt template, LLMs can leverage their instruction-following ability to provide reliable evaluation. For example, Li et al. (2023b) constructed a test set containing 805 questions and used the win rate compared with text-davinci-003 as the evaluation result, which is determined by GPT4. Zheng et al. (2023) developed 80 multi-round test questions covering eight common areas, and then automatically scored the model’s answers using GPT4. Their results reveal that strong LLM-based evaluators can achieve a high agreement rate among human experts, establishing a foundation for LLM-as-a-Judge.
@@
Model Foundation Instruction Response Annotation Evaluation Scheme Testset
@@
(Alpaca,Vicuna…) GPT4 Pairwise Grading GPT4
@@
(Kim et al., 2023) LLaMA2-chat GPT4 Generated GPT4 Generated GPT4 Pointwise Grading GPT4
Table 1: Detailed statistics of the four fine-tuned judge models, which is the foundation of our cross-validation. All the four models are open-source, with their training and test data also publicly released.
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. To address these issues, several fine-tuned judge models are proposed, relying on open-source foundation models and data constructed from either GPT4 or human annotation, as shown in Figure 1. For instance, PandaLM Wang et al. (2024) constructs data based on Alpaca instructions and GPT3.5 annotation, and then fine-tunes LLaMA-7B Touvron et al. (2023) as a judge model. JudgeLM Zhu et al. (2023b) constructs data from GPT4 annotations and fine-tunes a scalable judge model. Auto-J Li et al. (2023a) constructs judgement data upon multiple scenarios to train a generative judge model, which can provide both judgement and critic. Prometheus Kim et al. (2023) defines thousands of evaluation criteria and finetunes a fine-grained judge model.
In this paper, we conduct an empirical study of the evaluation capabilities of judge models. We conduct extrapolated evaluations among available judge models and meta-evaluation testsets. Experiment results indicate that while the fine-tuned judge models achieve superior accuracy on their respective in-domain test sets, they still exhibit limitations in the following aspects:
  • The fine-tuned judge model is inherently a classification model;
  • The fine-tuned judge model is overfitted to specific evaluation schemes;
  • The fine-tuned judge model is biased towards superficial quality;
To draw a conclusion, the fine-tuned judge models should be used only in similar evaluation scenarios, and can not serve as a general substitution for GPT4 in terms of LLM evaluation.
@@ 2 How Far can Fine-tuned Judges Go?
Step 1: Data Collection. The training data generally comprises three components: instructions, responses and evaluations. The instructions are typically obtained from instruction datasets, with the responses generated by various representative models, and the evaluations can be derived from either GPT4 or human annotation;
Step 2: Prompt Designing. The prompt template can be structured in various ways depending on the evaluation scheme, such as pairwise selection (which aims to select the better one from a pair of responses), pointwise grading (which aims to score a single reference), etc.
Step 3: Model Fine-tuning. Using the designed prompt and collected data, the training process of the judge model typically follows the instruction fine-tuning paradigm Ouyang et al. (2022). The model is fed with a instruction alongside answer(s) for yielding output comprising evaluation results.
While current judge models mostly self-claim exceeding the evaluation capability of GPT4, in this work, we make an cross validation based on four representative works as listed in Table 1. Our findings are presented in the following sections.
Model Train JudgeLM-test PandaLM-test Auto-J-test Prometheus-test
accuracy F1 accuracy F1 agreement pearson-ind pearson-ood
GPT 3.5 73.83 52.85 62.96 58.20 42.7 0.636 0.563
GPT 4-0613 85.28 76.87 78.68 73.24 56.3 0.742 0.743
[…]
2.1 Inherently a Classification Model
If we do not consider critic generation (which is seldom used in system-level evaluation), the LLM evaluation is inherently a classification (or regression) task. In contrast to the current judge models trained in a generation-style, we train four classification (regression) models based the four groups of data in Table 1111The training settings and prompt templates are presented in Appendix A.1 due to space limitation.. We also train four classification models based on DeBERTaV3-large He et al. (2023) on the same data, which is 20 times smaller than the 7B version of LLaMA.
As shown in Table 2, the classification model performs equally well as the generation model for both pairwise selection and pointwise grading. The powerful generation ability of LLMs hardly brings any improvement to the evaluation accuracy, as they are trained on the same data for the same objective. Moreover, the DeBERTa-based evaluator achieves comparable performance with the LLM-based evaluators [^2] The only exception is on Auto-J-test, which is possibly due to a large proportion of the test data exceeds 512 (the maximum context length of DeBERTa)., which might be argued for that the encoder-only architecture is more suitable for classification.
We also analyze the correlation between different predictions made by different evaluators. As shown in Figure 2 and 3, the correlation among different classification models is much closer than their correlation with GPT4. Different as they are in architectures, all three models are inherently classifiers fitting to the same set of supervision, leading to similar evaluation outcomes.
2.2 Overfitting to Evaluation Scheme
@@
GPT 3.5 73.83 52.85 62.96 58.20 42.7 59.83
GPT 4-0613 85.28 76.87 78.68 73.24 56.3 73.42
JudgeLM-7B 79.02 71.87 70.97 67.59 46.6 65.53
PandaLM-7B 65.24 47.42 67.57 57.49 40.0 57.61
Auto-J-13B 72.86 57.60 71.47 61.01 54.6 66.31
@@
One of the most appealing attribute of LLMs is their generalization ability, enabling them to execute various task defined by various instructions Zhu et al. (2023a). Under the case of LLM evaluation, the instruction can also be formed in various schemes: pairwise selection, pointwise grading, etc. Since different judge models are fine-tuned on different schemes, we could verify their evaluation capability under the scheme defined by others. Specifically, we cross-validate the judge models on each other’s testset [^3] We applying their publicly released checkpoints with predefined prompts. For details please refer to Appendix A.2..
As shown in Table 3 and 4, all four models perform the best on their own training settings, respectivly, with results comparable with GPT4. However, if we employ a model on evaluation schemes where it is not trained on [^4] For example, using a pairwise model (such as PandaLM or JudgeLM) for pointwise grading (such as Promethues testset), or using a pointwise model (such as Promethues) for pairwise selection (such as PandaLM or JudgeLM testsets)., the evaluation performance would drop by a large margin. On the contrary, GPT4 consistently exhibits superior performance across various evaluation schemes.
We also validate the judge models on MT-bench Zheng et al. (2023), which is a multi-turn meta-evaluation dataset. As can be seen in Table 5, while the four models are all trained for single-turn evaluation, they underperforms GPT4 on MT-bench by a large margin. This demonstrates that the finetuned judge models are overfitted to their respective evaluation schemes and lack generalibility.
@@ 2.3 Biased Towards Superficial Quality
Recently, there has been a few researches on the bias of LLM-based evaluators, namely the evaluator would favor more verbose answers, or answers with similar format Wang et al. (2023); Saito et al. (2023). To address this issue, Zeng et al. (2023) proposed LLMBar as a testbed for the fairness of evaluators. It comprises one natural testset (Natural) and four adversarial testsets (Neighbor, Manual, GPTOut, GPTInst), and the adversarial testsets consist of paired outputs with a correct answer and a smooth, coherent but inconsistent answer. We evaluate the judge models on LLMBar\@footnotemark, and the results are shown in Table 6.
It seems that the fine-tuned judge models are severely biased to superficial quality such as formality or verbosity, while neglecting crucial properties such as instruction following, resulting an accuracy even worse than random guess on the adversarial testsets. On the other hand, GPT4 does not over-rely on the superficial features and achieves decent accuracy on all the testsets. This also verifies that the fine-tuned judge models are inherently classifiers overfitted to the training data, relying on spurious statistical features for prediction.
It deserves noticing that the DeBERTa-based evaluator also outperforms the LLM-based evaluator by a large margin in terms of fairness. This inspires us that the bias of LLM-based evaluator may come from the casual language modeling process. While the model is trained to generate fluent and verbose responses, it also tends to prefer fluent and verbose response when employed for evaluation, even if it is not aligned with the instruction.
@@
Table 6: Accuracy of evaluators on bias evaluation.

### v2

# On the Limitations of Fine-tuned Judge Models for LLM Evaluation

@@ Abstract
Recently, there has been a growing trend of utilizing Large Language Model (LLM) to evaluate the quality of other LLMs. Many studies have employed proprietary close-source models, especially GPT-4, as the evaluator. Alternatively, other works have fine-tuned judge models based on open-source LLMs as the evaluator. While the fine-tuned judge models are claimed to achieve comparable evaluation capability with GPT-4, in this study, we conduct an empirical study of judge models. Our findings indicate that although the fine-tuned judge models achieve high performance on in-domain test sets, even surpassing GPT-4, they underperform GPT-4 across several dimensions, including generalizability, fairness, aspect-specific evaluation, and scalability. We also reveal that the fine-tuned judge model inherently operates as a task-specific classifier, consequently imposing the limitations. Finally, we propose an effective indicator to measure the reliability of fine-tuned judges, with the aim of maximizing their utility in LLM evaluation [^1] Codes are openly available at https://github.com/HuihuiChyan/UnlimitedJudge.
@@ 1 Introduction
Recently, the evaluation for Large-scale Language Models (LLMs) has drawn considerate attention of research community Liang et al. (2022); Chang et al. (2023). As the capabilities of LLMs continue to develop across various tasks, it is essential to evaluate them from comprehensive perspectives Qin et al. (2023). However, existing benchmarks, such as MMLU Hendrycks et al. (2021) and BIG-bench bench authors (2023), cannot fully showcase the generative ability of LLMs.
Figure 1: The general training and inference procedure of fine-tuned judge models. The training data generally comprises triplets with instruction, response and evaluation result, which are fed to the open-sourced foundation model (such as LLaMA) to create a judge model.
@@
Some research has proposed LLM-as-a-Judge Li et al. (2023c); Zheng et al. (2023), namley utilizing proprietary LLMs, especially GPT-4 Achiam et al. (2023), to evaluate the LLM’s response. By defining evaluation schemes in the prompt template, LLMs can leverage their instruction-following ability to provide reliable evaluation, achieving a high agreement rate with human evaluators.
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. To address these issues, several fine-tuned judge models are proposed Zhu et al. (2023b); Wang et al. (2024), relying on open-source foundation models and data constructed from either GPT-4 or human annotation, as shown in Figure 1. These models are validated on their respective meta-evaluation benchmarks, where the finetuned models exhibit performance on par with GPT-3.5 and GPT-4, leading to the affirmation of their evaluation capability.
In this paper, we conduct an empirical study of the evaluation capabilities of different judge models. Our study encompasses a comprehensive comparison across various benchmarks and dimensions. Experiment results indicate that while the fine-tuned judge models achieve superior accuracy on their respective in-domain test sets, they still exhibit limitations compared with close-sourced models in the following aspects:
  • The fine-tuned judge model is constrained by specific evaluation scheme;
  • The fine-tuned judge model is biased towards superficial quality;
  • The fine-tuned judge model is incapable of aspect-specific evaluation;
  • The fine-tuned judge model can not benefit from prompt engineering;
We think these limitations primarily stem from the fine-tuning process, where the foundation model is transformed into a task-specific classifier overfitted to the fine-tuning data.
Finally, we propose an indicator to quantify the reliability of the fine-tuned judge when applied for LLM evaluation. Our indicator is based on confidence estimation from softmax probability distribution, and we apply calibration to better model the confidence brought by the task-specific fine-tuning process. We validate our indicator for the judge model when applied to different meta-evaluation benchmarks, and managed to select the samples lying inside the scope of fine-tuning, on which the performance of the judge is more reliable.
Our contributions can be summarized as follows:
  • We reveal several limitations of fine-tuned judge models for LLM evaluation by conducting a comprehensive empirical study;
  • We propose an indicator to quantify the reliability of the fine-tuned judge, which can help to select reliable evaluations;
  • To the best of our knowledge, this is the first systematic study of the limitations of fine-tuned judge models for LLM evaluation;
2 Related Work
As Large Language Models (LLMs) continue to excel across various tasks and domains, it is essential to design an efficient and effective evaluation method. However, traditional evaluation metrics for generative models, such as BLEU Papineni et al. (2002) and ROUGE Lin (2004), are not expressive enough to recognize the good responses. On the other hand, after the emergence of BERT Kenton and Toutanova (2019), there are also several benchmarks designed to evaluation the capabilities of language models, such as MMLU Hendrycks et al. (2021) and BIG-bench bench authors (2023). But they are mostly formalized as multi-choice selection, which cannot not fully showcase the models’ generation capability.
Recent research has introduced the concept of LLM-as-a-Judge Li et al. (2023c); Zheng et al. (2023), namely utilizing proprietary LLMs, especially GPT4 Achiam et al. (2023), to evaluate the LLM’s response. For example, Li et al. (2023c) constructed a test set containing 805 questions and used the win rate compared with text-davinci-003 as the evaluation result, which is determined by GPT-4. Zheng et al. (2023) developed 80 multi-round test questions covering eight common areas, and then automatically scored the model’s answers using GPT-4. The GPT-4-based evaluator is proven to be high accurate compared with professional human evaluators, and presenting even better consistency and stability compared with human.
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. Therefore, follow-up works suggest fine-tuning language models specialized in evaluations. For instance, PandaLM Wang et al. (2024) constructs data based on Alpaca instructions and GPT-3.5 annotation, and then fine-tunes LLaMA-7B Touvron et al. (2023) as a judge model. JudgeLM Zhu et al. (2023b) constructs data from diversified instruction sets and GPT-4 annotations, and fine-tunes Vicuna Chiang et al. (2023) as a scalable judge model. Auto-J Li et al. (2023a) constructs judgement data upon multiple scenarios to train a generative judge model, which can provide both judgement and critic. Prometheus Kim et al. (2023) defines thousands of evaluation criteria and construct a feedback dataset based on GPT-4, and finetunes a fine-grained judge model.
Despite relying on models with 7B or 13B parameters, these fine-tuned judge models all achieve comparable accuracy with GPT-3.5 or GPT-4 on their respective testsets. However, the evaluation is mostly conducted on the in-domain testsets constructed similarly with the training data. A thorough examination to the evaluation capability of the fine-tuned judges is of urgent demand.
{tblr}
@@ 2 Related Work
hlines, Model & Foundation Instruction Response Annotation Evaluation Scheme Testset
@@
(Alpaca,Vicuna…) GPT-4 Pairwise Grading GPT-4
@@
(Kim et al., 2023) LLaMA2-chat GPT-4 Generated GPT-4 Generated GPT-4 Pointwise Grading GPT-4
Table 1: Detailed statistics of the four fine-tuned judge models, which is the foundation of our empirical study. All the four models are open-source, with their training and test data also publicly released.
@@ 3 How Far can Fine-tuned Judges Go?
In this section, we make a comprehensive empirical study based on four representative fine-tuned judge models as listed in Table 1. Section 3.1 offers a brief introduction of the construction of fine-tuned judges, and the subsequent sections explain their limitations one by one.
3.1 Preliminary: Fine-tuned LLM Judge
@@
Step 1: Data Collection. The training data generally comprises three components: instructions, responses and evaluations. The instructions are typically obtained from instruction datasets, with the responses generated by various representative models, and the evaluations can be derived from either GPT-4 or human annotation.
Step 2: Prompt Designing. The prompt template can be structured in various ways depending on the evaluation scheme, such as pairwise selection (which aims to select the better one from a pair of responses to the instruction), and pointwise grading (which aims to assign a score to a single response based on the instruction), etc.
Step 3: Model Fine-tuning. Using the designed prompt and collected data, the training process of the judge model typically follows the instruction fine-tuning paradigm Ouyang et al. (2022). The model is fed with a instruction alongside response(s) to generate output, which includes evaluation results and possibly explanations.
After that, the fine-tuned judge can be adopted in evaluating the output of LLM. While these models are able to achieve superior performance on their self-designed testsets, we reveal that there exists several limitations about the evaluation capabilities of the fine-tuned judges.
@@
JudgeLM-7B 78.98 68.62 68.17 64.87 46.6 64.58
PandaLM-7B 66.44 56.01 68.97 60.95 40.0 58.47
Auto-J-13B 77.19 60.42 72.27 64.27 54.6 68.02
@@ 3.2 Constrained by Evaluation Scheme
One of the most appealing attribute of LLMs is their generalization ability, enabling them to execute various task defined by various instructions Zhu et al. (2023a). Under the case of LLM evaluation, the instruction can also be formed in various schemes: pairwise selection, pointwise grading, etc. Since different judge models are fine-tuned on different schemes, we would like to verify their evaluation capability under the scheme defined by others. Specifically, we apply their publicly released checkpoints, and cross-validate the judge models on each other’s testsets [^2] We make minimal modification to the predefined prompts to adapt the model to different schemes. For detailed prompts please refer to Appendix A.2..
As shown in Table 2 and 3, all four models perform the best on their own training schemes, respectively, with results comparable with GPT-4. However, if we employ a model on an evaluation scheme where it is not trained on, the evaluation performance would drop by a large margin. For example, using a pairwise model (such as PandaLM or JudgeLM) for pointwise grading (such as Prometheus testset), or using a pointwise model (such as Prometheus) for pairwise selection (such as PandaLM or JudgeLM testsets), would all lead to catastrophic performance degradation. On the contrary, close-sourced models such as GPT-3.5 or GPT-4 consistently exhibits superior performance across various evaluation schemes.
We also validate the judge models on MT-bench Zheng et al. (2023), which is a multi-turn meta-evaluation dataset\@footnotemark. As shown in Table 4, while the four models are all trained for single-turn evaluation, they underperforms GPT-4 on MT-bench by a large margin. This demonstrates that the finetuned judge models are overfitted to their respective evaluation schemes and lost their generalibility.
@@ 3.3 Biased Towards Superficial Quality
Recently, there has been a lot research on the bias of LLM-based evaluators, namely the evaluator would favor more verbose answers, or answers with similar format Wang et al. (2023b); Saito et al. (2023). To address this issue, Zeng et al. (2023) proposed LLMBar as a testbed for the fairness of evaluators. It comprises one natural testset (Natural) and four adversarial testsets (Neighbor, Manual, GPTOut, GPTInst), and the adversarial testsets consist of paired outputs with a correct answer and an incorrect answer with better superficial quality (e.g., more fluent, more verbose, etc.). We evaluate the judge models on LLMBar [^3] The detailed prompts are present in Appendix A.2. as shown in Table 5.
@@
GPT-4-1106 93.5 64.2 76.6 76.6 75.0
Table 5: Accuracy of evaluators on bias evaluation. Neig., GPTI., GPTO., and Manu. are the four adversarial test sets designed to quantify the bias.
As can be seen, the fine-tuned judge models achieves a poor results on adversarial testsets, even much worse than random-guess. This notifies that they are severely biased to superficial quality such as formality or verbosity, while neglecting crucial properties such as instruction following, resulting in the preference to the incorrect answers. On the other hand, GPT-4 does not over-rely on the superficial features and achieves decent accuracy on all the testsets. This notifies that the superior performance of fine-tuned judges on the in-domain testsets may rely on spurious statistical features Niven and Kao (2019), instead of really differentiating good and bad responses.
HaluEval-QA HaluEval-Sum HaluEval-Dial ToxicChat SALAD-Bench
Model accuracy F1 accuracy F1 accuracy F1 accuracy F1 accuracy F1
JudgeLM-7B - - - - - - - - 82.45 57.44
PandaLM-B - - - - - - - - 57.03 37.23
Auto-J-13B 58.30 56.03 53.10 43.34 63.10 62.90 87.40 52.24 86.88 52.66
w/o adapt 59.60 57.38 53.47 43.55 64.50 63.71 87.70 51.15 71.77 47.86
Prometheus-7B 47.90 45.84 44.50 40.38 51.00 45.17 77.10 58.14 - -
w/o adapt 48.90 45.10 46.60 36.43 53.40 50.24 81.20 61.87 - -
GPT-3.5-0613 57.50 57.10 62.60 60.27 72.10 72.08 95.10 80.80 98.75 97.54
GPT-4-1106 72.50 72.50 72.00 71.44 84.50 84.78 94.50 82.78 100 100
Table 6: Results of evaluators on aspect-specific evaluation. w/o adapt denotes using the original prompt without adaptation to the specific aspect. As HaluEval and ToxicChat are both binary classification, we apply Auto-J and Prometheus with pointwise grading and conduct grid-search to determine the classification threshold. On the other hand, as SALAD-Bench is pairwise classification, we apply pairwise selection models, namely JudgeLM, PandaLM and Auto-J to select a better response.
We also fine-tune a DeBERTa-based judge model in a classification style on LLMBar (please refer to Section 3.6 for details). It deserves noticing that the DeBERTa-based evaluator also outperforms the LLM-based evaluator by a large margin in terms of fairness. This inspires us that the bias of LLM-based evaluator may come from the casual language modeling process. While the model is trained to generate fluent and verbose responses, it also tends to prefer fluent and verbose response when employed for evaluation, even if it is not aligned with the instruction.
3.4 Incapable of Aspect-specific Evaluation
LLM evaluation covers various aspects such as usefulness, safety, and factuality, and sometimes we are particularly interested in a specific aspect. While previous work primarily assess the evaluation capability of the judge models from a general perspective, we would like to assess them on fine-grained aspects. We select the following three datasets:
  1. 1.
    HaluEval Li et al. (2023b): This dataset focuses on factuality evaluation. It contains generated and human-annotated hallucinated samples, which lie in three domains: Question-Answering, Summary and Dialogue. Given an instruction-response pair, the evaluator should decide whether the response is hallucinated.
  2. 2.
    ToxicChat Lin et al. (2023): This dataset focuses on toxicity evaluation. It contains toxic and non-toxic conversations based on real-world user-AI interactions. Given an instruction-response pair, the evaluator should decide whether the response is toxic.
  3. 3.
    SALAD-Bench Li et al. (2024): This dataset focuses on safety evaluation. It contains instructions and responses spanning among different domains and tasks. Given an instruction and a pair of responses, the evaluator should decide which response is safer.
We validate both close-sourced and fine-tuned judges on the three datasets [^4] We make minimal modification to the prompts to adapt them to the specific aspects, as detailed in Appendix A.2.. As can be seen from Table 6, the fine-tuned judges fall far behind the close-sourced judges on all fine-grained aspects. It deserves to notice that while Prometheus is designed for fine-grained evaluation, it obtains an inferior performance on both benchmarks, which notifies that it failed to learn the correlation between fine-grained aspects and evaluation results.
For the purpose of comparison, we also apply Auto-J and Prometheus with their original prompt on aspect-specific evaluation. As can be seen in Table 6, to our surprise, their performance remains roughly the same compared with aspect-specific prompts, notifying that both models have lost the general instruction-understanding ability, therefore the aspect-specific prompt is not taking effect.
3.5 No Benefits from Prompt Engineering
One of the most appealing features of LLM is it can benefit from delicate prompt engineering. Various strategies have been proposed to improve the LLM’s capability on various tasks, including text evaluation. In this section, we select two representative strategies, namely In-context Learning (ICL) Dong et al. (2023) and Chain-of-Thought Prompting (CoT) Wei et al. (2022), to further improve the evaluation capability of the judge models:
JudgeLM-test PandaLM-test
Model accuracy F1 accuracy F1
JudgeLM-7B 78.98 68.62 68.17 64.87
+ CoT 77.68 67.59 68.03 64.42
+ ICL 68.57 58.52 41.14 40.39
PandaLM-7B 66.44 56.01 68.97 60.95
+ CoT 65.85 56.59 68.03 60.42
+ ICL 66.16 55.94 68.97 59.40
Auto-J-13B 77.19 60.42 72.27 64.27
+ ICL 76.20 59.12 68.37 58.44
GPT-3.5-0613 72.57 51.40 64.36 46.40
+ CoT 75.24 60.71 69.97 63.66
+ ICL 69.38 57.46 70.67 56.12
GPT-3.5-0125 70.67 50.44 64.46 46.60
+ CoT 69.24 56.51 73.37 65.89
+ ICL 70.24 60.46 70.37 53.65
GPT-4-1106 85.28 76.87 74.07 68.09
+ CoT - - 77.08 71.77
+ ICL - - 64.86 56.20
Table 7: Results of evaluators with ICL and CoT. Increased results are in bold while decreased results are in grey. We did not apply GPT-4 on JudgeLM-test as the annotation of JudgeLM-test is conducted with GPT-4 without ICL and CoT. We only apply ICL on Auto-J as the original prompt of Auto-J comprises CoT.
  1. 1.
    In-context Learning (ICL): where task demonstrations are integrated into the prompt as illustration. In our work, we randomly select 2-4 ICL demonstrations from the training set based on the max-context length.
  2. 2.
    Chain-of-thought (CoT): where the input prompt is structured in a way that mimics human reasoning. In our work, the judge model is forced to generate the explanation first, then provide a final judgement.
We validate both close-sourced and fine-tuned judges with the two strategies [^5] The detailed prompts are presented in Appendix A.2.. As shown in Table 7, the close-sourced models are improved by a large margin through both prompt engineering strategies. Conversely, the fine-tuned judges hardly benefit from these strategies, sometimes even experiencing severe performance decline. Specifically, in the case of CoT prompting, despite we modified the prompts for JudgeLM and PandaLM to generate CoT firstly, both models adhered to their original output format and failed to produce CoT. While there exist more intricate prompting strategies such as ChatEval Chan et al. (2024) or Branch-Solve-Merge Saha et al. (2023), we posit they can neither bring benefit to the fine-tuned judges, as they have lost their general instruction-following ability and are constrained to a singular output pattern.
3.6 The Essence of Fine-tuned Judge: A Task-specific Classifier
Model JudgeLM-test PandaLM-test Auto-J-test Prometheus-test
accuracy F1 accuracy F1 agreement pearson-ind pearson-ood
Released Models 78.98 68.62 68.97 60.95 54.6 0.864 0.869
Vicuna-generation 82.44 71.77 72.37 60.78 47.6 0.826 0.815
Vicuna-classification 82.16 70.07 70.87 60.34 46.8 0.846 0.831
DeBERTa-classification 81.30 68.34 72.27 51.75 31.7 0.835 0.813
GPT-3.5-0613 72.57 51.40 64.36 46.40 42.7 0.636 0.563
GPT-4-1106-preview 85.28 76.87 74.07 68.09 56.3 0.742 0.743
Table 8: Comparison of generation and classification-based evaluators. Results with are from evaluating the four publicly released models on their respective testsets, and results with are from evaluating models trained by us.
Combining all the limitations revealed in our experiments, we would like to claim that after the fine-tuning process on a single task, the judge model has degenerated into a task-specific classifier, which is overfitted to the training data. To support this, we fine-tune three groups of judges based on the four groups of data as listed in Table 1666Please refer to Appendix A.1 for training details.:
  1. 1.
    Vicuna-generation: It formulates the evaluation task in a generation-style, and the prediction head reuses the pretrained language model head, and is trained akin to the process of language modeling, based on the 7B version of Vicuna Chiang et al. (2023);
  2. 2.
    Vicuna-classification: It formulates the evaluation task as classification or regression, and the prediction head is newly initialized as a linear projection layer, and is decoupled from the language modeling process;
  3. 3.
    DeBERTa-classification: It also formulates as a classification task, based on DeBERTaV3-large He et al. (2023), which is 20 times smaller than the 7B version of Vicuna;
Notice for fine-tuning Vicuna-generation and Vicuna-classification models, we adopt the same prompt and hyper-parameters, with the only difference lying in the prediction method [^7] An illustration figure is presented in Appendix A.1..
As shown in Table 8, the classification model performs equally well as the generation model on both pairwise selection and pointwise grading. The formidable generative capabilities of LLMs hardly bring any improvement to the evaluation, as they are fitting to the same group of data. Moreover, the DeBERTa-based classifier achieves comparable performance with the LLM-based evaluators [^8] The only exception is on Auto-J-test, which is possibly due to a large proportion of the test data exceeds 512 (the maximum context length of DeBERTa)., which might be argued for that the encoder-only architecture is more suitable for classification.
We also analyze the correlation between different predictions made by different evaluators. As shown in Figure 2 and 3, the correlation among different classification models is much closer than their correlation with GPT-4. Different as they are in architectures, all three models are inherently classifiers fitting to the same set of supervision, leading to similar evaluation outcomes.
Although prior research on instruction-tuning all emphasizes the importance of data diversity Zhou et al. (2023); Lu et al. (2024), the fine-tuning of LLM judges is doing the opposite thing. Therefore, after fine-tuning for a single task with a fixed prompt template, the model lost its generalization ability, and degenerate into a task-specific classifier, which exhibits several limitations due to overfitting.
4 When to Trust Fine-tuned Judges?
Model Method SALAD-Bench JudgeLM-test PandLM-test Auto-J-test Average
accuracy F1 accuracy F1 accuracy F1 agreement
JudgeLM-7B random 82.50 57.36 80.40 71.87 67.54 65.20 43.25 59.42
perplexity 88.09 59.57 85.48 74.38 72.34 63.25 49.57 61.69
SoftEnt 95.10 63.62 89.40 78.09 77.96 64.01 55.89 65.40
SoftEnt-Cali 91.56 61.47 88.70 76.30 79.76 67.05 56.32 65.29
PandaLM-7B random 56.67 36.67 68.09 48.26 66.73 58.86 39.37 45.79
perplexity 59.37 55.03 70.15 52.62 68.53 57.73 42.53 51.98
SoftEnt 63.13 62.07 72.96 53.67 75.55 62.53 43.53 55.45
SoftEnt-Cali 66.77 63.77 73.91 55.97 76.35 67.30 45.40 58.11
Auto-J-13B random 72.29 48.10 77.27 60.7 71.54 63.51 46.12 54.61
perplexity 77.47 52.56 80.28 61.85 75.55 65.84 50.35 57.65
SoftEnt 81.98 54.68 79.13 63.58 80.16 66.61 53.01 59.47
SoftEnt-Cali 82.08 54.75 80.19 63.58 81.36 70.00 52.30 60.16
Table 9: Comparison of different reliability indicators for the judge models. We split the test sets into halves based on different indicators, and report the performance of the judge on the half with higher scores. The on random-split baseline denotes that the results are averaged over three times.
Despite the limitations revealed in our study, we aim not to disregard the significance of fine-tuned judges entirely. While LLM exhibit excellent performance among various tasks, task-specific models are still wildly used everywhere in natural language processing. Therefore, it deserves our discussion about the reliability of the fine-tuned judges. For this purpose, we propose a quantitative indicator to estimate whether an sample can be reliably evaluated by a judge model.
Borrowing the idea from confidence estimation Huang et al. (2024), we propose to quantify the reliability of the judge based on softmax entropy. Given an instruction $x$ and a fine-tuned judge model with parameters $\theta$, the reliability of generating response $y$ can be factorized as:
$\textrm{SoftEnt}(y|x,\theta)=-\frac{1}{T}\sum_{t=1}^{T}\sum_{v=1}^{V}p_{\theta}(y_{t}^{v})\mathrm{log}p_{\theta}(y_{t}^{v})$
where $p(y_{t})$ represents the conditional distribution $p(y_{t}|x,y_{<t},\theta)$ at each decoding step, $T$ is the response length, and $V$ is the vocabulary size.
As we would like to quantify whether the sample lies in the task-specific fine-tuning scope, we further calibrate the SoftEnt as follows:
[…]
where $\theta_{j}$ denotes the fine-tuned judge model, and $\theta_{b}$ denotes its corresponding foundation model. By calibration, we aim to exclude the influence of foundation model, thus modeling solely the confidence instilled by the task-specific fine-tuning process.
(a) JudgeLM-7B applied on SALAD-Bench
(b) Auto-J-13B applied on SALAD-Bench
Figure 4: Accuracy of judge models when applied to different buckets of data grouped with reliability scores.
To verify the effectiveness of the reliability score, we re-conduct the cross-validation in Section 3, wherein the test set was split into halves based on different indicators. As shown in Table 9, our proposed SoftEnt-Cali managed to select the samples with better reliability, therefore achieves a higher accuracy than random split. On the other hand, split based solely on the SoftEnt or perplexity of the judge model underperforms, as they did not exclude the influence of the foundation model.
We conducted an additional experiment wherein samples were grouped into buckets based on their SoftEnt-Cali scores, allowing us to assess the judge model across different buckets. As shown in Figure 4, the accuracy of judge models exhibit a strong correlation with the SoftEnt-Cali score. Notably, for both judges on SALAD-bench, buckets with higher scores demonstrated an accuracy of 90%, whereas those with lower scores performed akin to random guessing. This underscores the efficacy of SoftEnt-Cali as a reliability indicator of the judge model’s performance for evaluation.
The reliability indicator serves as a useful tool to compensate for the limitations and improve the LLM evaluation. For example, it can be used to decide whether an evaluation sample requires human intervention, or to select the most reliable one when multiple judge models are available.